Iam having a horrible time w/quadratic equations especially finding decimals here is the first one: duringthe first part of a trip a canoeist travels 91 mi. at a certain speed. The canoeist travels 15 mi on the second part of the trip at a speed 5 mph slower. The total time for the trip is 4 hours. What was the speed on each part of the trip. The hardest time I am having is to use the quadratic formula and round to the nearest hundreth PLEASE help me understand this. How do you divide quadratic equations to get decimals??
Answer by stanbon(75887)

duringthe first part of a trip a canoeist travels 91 mi. at a certain speed. The canoeist travels 15 mi on the second part of the trip at a speed 5 mph slower. The total time for the trip is 4 hours. What was the speed on each part of the trip.
---
1st Part DATA:
distance = 91 mi ; rate = x mph ; time = d/r = 91/x hrs.
---
2nd Part DATA:
distance = 15 mi ; rate = (x-5) mph ; time = d/r = 15/(x-5) hrs.
---
Equation:
time + time = 4 hrs
91/x + 15/(x-5) = 4
Mltiply thru by x(x-5) to get:
91(x-5) + 15x = 4x(x-5)
91x - 455 + 15x = 4x^2 - 20x
4x^2 - 126x + 455 = 0
---
Quadratic Formula:
x = [126 +- sqrt(126^2 - 4*4*455)]/(2*4)
---
x = [126 +- sqrt(8596)]/8
---
x = [126 +- 92.71]/8
----
x = [126+92.71]/8 or x = [126-92.71]/8
---
x = 27.34 or x = 4.16
===============================
Cheers,
Stan H.